ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 247, 2022

Alesker, S. Kotrbatý’s theorem on valuations and geometric inequalities for convex bodies

361

Alves, C. O. and J. A. dos Santos Multiple solutions for semilinear discontinuous variational problems with lack of compactness

149

Antony, A. E., P. Komma and V. Z. Thomas A property of pgroups of nilpotency class p + 1 related to a theorem of Schur

251

Avilés, A., G. Martínez-Cervantes, J. Rodríguez and P. Tradacete A Godefroy–Kalton principle for free Banach lattices

433

Bleher, F. M., T. Chinburg and J. Gillibert Unramified Heisenberg group extensions of number fields

233

Chatzidakis, Z. and P. Zalesskii Pro-p groups acting on trees with finitely many maximal vertex stabilizers up to conjugation

593

Chinburg, T. See F. M. Bleher Colarte-Gómez, L., E. Mezzetti, R. M. Miró-Roig and M.

Salat-Moltó Togliatti systems associated to the dihedral group and the weak Lefschetz property

195

Damanik, D. and D. Lenz Absence of absolutely continuous spectrum for generic quasi-periodic Schrödinger operators on the real line

783

Dolgopyat, D., C. Dong, A. Kanigowski and P. Nándori Mixing properties of generalized T, T^{−1} transformations

21

Dong, C. See D. Dolgopyat Dostert, M., A. Kolpakov and F. M. Oliveira Filho Semidef-

inite programming bounds for the average kissing number

635

and P. Shumyatsky On profinite groups with automorphisms whose fixed points have countable Engel sinks

303

Finkelberg, M., D. Kazhdan and Y. Varshavsky Lusztig conjectures on S-cells in affine Weyl groups

379

Fuhrmann, G., M. Gröger and D. Lenz The structure of mean equicontinuous group actions

75

Gekhtman, D. and M. Greenfield Isometric submersions of Teichmüller spaces are forgetful

499

Geroldinger, A. and Q. Zhong A realization result for systems of sets of lengths 177 Ghara, S., S. Kumar and P. Pramanick 𝕂-homogeneous tuple of

operators on bounded symmetric domains

331

Gillibert, J. See F. M. Bleher Ginosar, Y. Realization-obstruction exact sequences for Clifford sys-

tem extensions

955

Glück, J. and F. G. Martin Uniform convergence of stochastic semigroups

1

Greenfield, M. See D. Gekhtman Gröger, M. See G. Fuhrmann Gurel-Gurevich, O. and M. Seidel Recurrence of a weighted ran-

dom walk on a circle packing with parabolic carrier

547

Han, J.-Z. Bimodules and universal enveloping algebras associated to VOAs

905

Herfort, W. and W. Hojka Archipelago groups are locally free,

Corrigendum to: “Cotorsion and wild homology”

993

Hirayama, M., D. H. Kim and Y. Son On the multiple recurrence properties for disjoint systems

405

Hojka, W. See W. Herfort

Huber, F. and C. Procesi Tensor polynomial identities  125
Johnson, W. Counting mod n in pseudofinite fields  697

Kanigowski, A. See D. Dolgopyat

Kato, F. Integral morphisms and log blow-ups  831

Kaufmann, U., H. Ramos Quoirin and K. Umezu Nonnegative solutions of an indefinite sublinear Robin problem II: local and global exactness results

661

Kawasaki, M. and M. Kimura Ĝ-invariant quasimorphisms and symplectic geometry of surfaces

845

Kazhdan, D. See M. Finkelberg Kimura, M. See M. Kawasaki Kim, D. H. See M. Hirayama Kochloukova, D. H. and C. Martínez-Pérez Coabelian ideals in

ℕ-graded Lie algebras and applications to right angled Artin Lie algebras

797

Kolpakov, A. See M. Dostert Komma, P. See A. E. Antony Kumar, S. See S. Ghara Lenz, D. See D. Damanik Lenz, D. See G. Fuhrmann Macedo, R., D. Pellegrino and J. Santos Nonlinear variants of

a theorem of Kwapień 217 Marcus, A. W., D. A. Spielman and N. Srivastava Interlacing

families III: Sharper restricted invertibility estimates

519

Martínez-Cervantes, G. See A. Avilés Martínez-Pérez, C. See D. H. Kochloukova Martin, F. G. See J. Glück Melbourne, I., N. Paviato and D. Terhesiu Nonexistence of

spectral gaps in Hölder spaces for continuous time dynamical systems

987

Mezzetti, E. See L. Colarte-Gómez Miró-Roig, R. M. See L. Colarte-Gómez

Mj, M. and P. Sardar Propagating quasiconvexity from fibersi  923

Nándori, P. See D. Dolgopyat Paviato, N. See I. Melbourne Pellegrino, D. See R. Macedo Pramanick, P. See S. Ghara Procesi, C. See F. Huber Ramos Quoirin, H. See U. Kaufmann Rodríguez, J. See A. Avilés Salat-Moltó, M. See L. Colarte-Gómez Santos, J. See R. Macedo dos Santos, J. A. See C. O. Alves Sardar, P. See M. Mj Schaeffer Fry, A. A. Galois-equivariant McKay bijections for

primes dividing q − 1

269

Seidel, M. See O. Gurel-Gurevich Shumyatsky, P. See Khukhro, E. I. Son, Y. See M. Hirayama Spielman, D. A. See A. W. Marcus Srivastava, N. See A. W. Marcus Tamam, N. Existence of non-obvious divergent trajectories in homo-

geneous spaces

459

Terhesiu, D. See I. Melbourne Thomas, V. Z. See A. E. Antony Tradacete, P. See A. Avilés Umezu, K. See U. Kaufmann

Vaccaro, A. Trivial endomorphisms of the Calkin algebra  873

Varshavsky, Y. See M. Finkelberg Velicu, A. and N. Yessirkegenov Rellich, Gagliardo–Nirenberg,

Trudinger and Caffarelli–Kohn–Nirenberg inequalities for Dunkl operators and applications

741

Yessirkegenov, N. See A. Velicu Zalesskii, P. See Z. Chatzidakis

Zheng, Z. On abelian automorphism groups of hypersurfaces  479

Zhong, Q. See A. Geroldinger de Oliveira Filho, F. M. See M. Dostert